Please help me solve this problem:
Adams Company bought a piece of land in 1981 for $200,000. By 2005, its value had increased to $1 million. Find the annual rate of appreciation during this period.

Answer by stanbon(75887)

Adams Company bought a piece of land in 1981 for $200,000. By 2005, its value had increased to $1 million. Find the annual rate of appreciation during this period.
----------------------------
Form: y = ab^x
---
You have two points: (0,200,000) and (24,1,000,000)
where the 1st coordinate is the number of years after 1981.
----------------------
Using (0,200,000 you get:
200,000 = ab^0
200,000 = a
------------------------
Using (24,1000000)
1000000 = 200,000b^24
5 = b^24
b = 5^(1/24) = 1.0694
---
Equation:
y = 200,000(1.0694)^x
----
Annual rate is 0.0694 or 6.94%
====================================
Cheers,
Stan H.
===============